A box contains 10 red and 12 white rose flowers. Four flowers are picked up at random one by one without replacement. What is the probability that;
a. The first 3 flowers are red?
b. three are 2 red and 2 white flowers
c. at least 2 are red.
1
Expert's answer
2020-11-30T13:06:08-0500

a. P(the first flower is red) = "\frac{10}{10 + 12} = \frac{10}{22}"

P(the probability for the second flower to be red, if the first flower is red) = "\frac{10-1}{9 + 12} = \frac{9}{21}"

P(the third flower is red, if the first and the second flowers are red) = "\frac{8}{20}"

P(the first 3 flowers are red) "= \frac{10}{22} \times \frac{9}{21} \times \frac{8}{20} = \frac{720}{9240} = 0.078" or 7.8 %

b. P(2 red and 2 white flowers) "= 6 \times \frac{10}{22} \times \frac{9}{21} \times \frac{12}{20} \times \frac{11}{19} = \frac{71280}{175560} = 0.4060" or 40 %

c. At least 2 are red out of 4. The number of red roses can be 2, 3 or 4.

Total number of red roses are 10.

If 2 roses are red:

10C2 "\times" 12C2 = "\frac{10!}{8!2!} \times \frac{12!}{10!2!} = 2970"

If 3 roses are red:

10C3 "\times" 12C1 = "\frac{10!}{7!3!} \times \frac{12!}{11!} = 1440"

If 4 roses are red:

10C4 "\times" 12C0 = "\frac{10!}{6!4!} \times 1 = 210"

P(at least 2 are red) "= \frac{2970 + 1440 +210}{C^{22}_4} = \frac{4620}{7315} = 0.631" or 63 %